# Defining an universal “sign” to strictly probe phase transition

### v3

@@ Abstract
Introduction.– In a quantum many-body system, the Hilbert space grows exponentially as the system size increases, and quantum Monte Carlo (QMC) tries to overcome the exponential-wall difficulty through importance sampling and focuses on a fraction of Hilbert space. Compared with other extensively used computing methods, such as tensor network or density matrix renormalization group, QMC is especially efficient in studying models in high dimensions [1, 2, 3, 4, 5, 6, 7].
@@
General analysis about sign.—
@@
Figure 3: (a) The free energy $F$ and $F_{+}$, as well as the their difference $\Delta F$ versus $T$. (b) The sign value as a function of $T$. (c) $\partial\langle sign\rangle/\partial T$ vs $T$. All the data are from ED calculations of the Eq.General analysis about sign.— and Eq.General analysis about sign.— on a 4$\times$4 square lattice with $J_{1}=1,g_{1}=0.7$ and $J_{2}=0.2,g_{2}=0.7$.
Negative examples Still, the failure of $\langle\mathrm{sign}\rangle$ in detecting PTP can be questioned as to whether the situation in Fig.2 (a2) and (a3) truly exists or not. This section provides practical numerical results in specific systems. Among several systems we examined, the anisotropic $J_{1}$-$J_{2}$ Heisenberg model on the square lattice can be a typical example. The $J_{1}$-$J_{2}$ Hamiltonian reads
@@
As displayed in Fig.3, the free energy $F$ of the system with Hamiltonian $H$ has no obvious mutation, while there is one in the curve of the free energy $F_{+}$ described by $H_{+}$, which is qualitatively the same case discussed in Fig.2(a2). One can extract a possible phase transition point around $T=0.7$ from the obvious change of $\langle\mathrm{sign}\rangle$ in Fig. 3(b) and the peak of its differential (c), which, however, only mistakenly refers to the possible critical point of the reference system $H_{+}$. We want to point out that although it is very hard to distinguish a singularity and a continuous peak in finite size, it should be noticed that either of them can be introduced by $H_{+}$, which would mix up the signal of the original system   [^1] According to the Mermin-Wagner theorem, we know this system has no finite temperature phase transition in fact. However, it is not important in our work because we are focusing on demonstrating the signal from the reference system can induce the sign value displaying a fake semaphore containing both a singularity (real phase transition) or a continuous peak (fake phase transition).. In this example, changing $\langle\mathrm{sign}\rangle$ would bring us confusing and misleading phase transitions. Similar results are also found on the spinful triangular Hubbard model where the onset of $\langle\mathrm{sign}\rangle^{\prime}$ starts away from the best-known location of the phase transition point [56].
Modified sign to probe the transition.– The conclusion from the previous analysis is slightly disappointing, but the successful situations inspire us. The original sign is defined as $\langle sign=Z(\beta,J)/Z_{+}(\beta,J)\rangle$ with $\beta=1/T$ and related to both systems, which is the decisive reason for its failure. If the denominator is fixed, the sign will behave like the $Z$ only, which can strictly reflect the phase transition. Therefore, a new kind of ”sign” can be redefined as
@@
Figure 4: The temperature dependence of $\partial^{2}[T\ln(\langle\widetilde{sign}\rangle)]/\partial T^{2}$ per site using SSE on the Hamiltonian in Eq.(General analysis about sign.—) from $L=4$ to $L=32$ on square lattice with $J_{1}=1$ and $g_{1}=g_{2}=-2$ The fixed calculation parameters of the reference system when calculating $\langle\widetilde{sign}\rangle$ are: $\beta^{*}=2/3$, $J_{2}^{*}=0.2$. The computing results at different temperatures for $L=4$ agree with the second-order derivative of free energy per site of the same Hamiltonian got from ED.
@@
As an example, we calculate $\langle\widetilde{sign}\rangle$ on one frustrated model in Eq.(General analysis about sign.—) with $g_{1}=g_{2}<-1$ using SSE. The interactions of spins in the z-direction are ferromagnetic with absolute values larger than the x- and y-direction, so there should be an Ising-like phase transition at finite temperature in this system. The second-order derivative of free energy got from $\langle\widetilde{sign}\rangle$ presents a diverging peak with size increases as shown in Fig.4, which unbiasedly probes the second-order phase transition point of the model.
Conclusions.—
@@
Acknowledgments.––We acknowledge Xu Zhang, Zi Yang Meng, Kai Sun for useful discussions. This research was supported by the National Natural Science Foundation of China (grant nos. 12004020, 12174167, 12247101), and the Fundamental Research Funds for the Central Universities in China. The authors acknowledge Beijng PARATERA Tech CO.,Ltd. for providing HPC resources that have contributed to the research results reported within this paper. ZY thanks the start-up fund of the Westlake University.

### v4

@@ Abstract
I Introduction
In a quantum many-body system, the Hilbert space grows exponentially as the system size increases, and quantum Monte Carlo (QMC) tries to overcome the exponential-wall difficulty through importance sampling and focuses on a fraction of Hilbert space. Compared with other extensively used computing methods, such as tensor network or density matrix renormalization group, QMC is especially efficient in studying models in high dimensions [1, 2, 3, 4, 5, 6, 7].
@@ II Positive examples by QMC
III General analysis about sign
@@
Figure 3: (a) The free energy $F$ and $F_{+}$, as well as the their difference $\Delta F$ versus $T$. (b) The sign value as a function of $T$. (c) $\partial\langle sign\rangle/\partial T$ vs $T$. All the data are from ED calculations of the Eq.IV and Eq.IV on a 4$\times$4 square lattice with $J_{1}=1,g_{1}=0.7$ and $J_{2}=0.2,g_{2}=0.7$.
IV Negative examples
Still, the failure of $\langle\mathrm{sign}\rangle$ in detecting PTP can be questioned as to whether the situation in Fig.2 (a2) and (a3) truly exists or not. This section provides practical numerical results in specific systems. Among several systems we examined, the anisotropic $J_{1}$-$J_{2}$ Heisenberg model on the square lattice can be a typical example. The $J_{1}$-$J_{2}$ Hamiltonian reads
@@ IV Negative examples
As displayed in Fig.3, the free energy $F$ of the system with Hamiltonian $H$ has no obvious mutation, while there is one in the curve of the free energy $F_{+}$ described by $H_{+}$, which is qualitatively the same case discussed in Fig.2(a2). One can extract a possible phase transition point around $T=0.7$ from the obvious change of $\langle\mathrm{sign}\rangle$ in Fig. 3(b) and the peak of its differential (c), which, however, only mistakenly refers to the possible critical point of the reference system $H_{+}$. We want to point out that although it is very hard to distinguish a singularity and a continuous peak in finite size, it should be noticed that either can be introduced by $H_{+}$, which would mix up the signal of the original system. Note: According to the Mermin-Wagner theorem, we know this system has no finite temperature phase transition in fact. However, it is not important in our work because we are focusing on demonstrating the signal from the reference system can induce the sign value displaying a fake semaphore containing both a singularity (real phase transition) or a continuous peak (fake phase transition). In this example, changing $\langle\mathrm{sign}\rangle$ would bring us confusing and misleading phase transitions. Similar results are also found in some Fermion models recently [55, 56]. The Ref.[56] also found the reference system can affect the sign with a fake phase transition.
V Modified sign to probe the transition
The conclusion from the previous analysis is slightly disappointing, but the successful situations inspire us. The original sign is defined as $\langle sign=Z(\beta,J)/Z_{+}(\beta,J)\rangle$ with $\beta=1/T$ and related to both systems, which is the decisive reason for its failure. If the denominator is fixed, the sign will behave like the $Z$ only, which can strictly reflect the phase transition. Therefore, a new kind of ”sign” can be redefined as
@@ V Modified sign to probe the transition
Figure 4: The temperature dependence of $\partial^{2}[T\ln(\langle\widetilde{sign}\rangle)]/\partial T^{2}$ per site using SSE on the Hamiltonian in Eq.(IV) from $L=4$ to $L=32$ on square lattice with $J_{1}=1$ and $g_{1}=g_{2}=-2$ The fixed calculation parameters of the reference system when calculating $\langle\widetilde{sign}\rangle$ are: $\beta^{*}=2/3$, $J_{2}^{*}=0.2$. The computing results at different temperatures for $L=4$ agree with the second-order derivative of free energy per site of the same Hamiltonian got from ED.
@@
As an example, we calculate $\langle\widetilde{sign}\rangle$ on one frustrated model in Eq.(IV) with $g_{1}=g_{2}<-1$ using SSE. The interactions of spins in the z-direction are ferromagnetic with absolute values larger than the x- and y-direction, so there should be an Ising-like phase transition at finite temperature in this system. The second-order derivative of free energy got from $\langle\widetilde{sign}\rangle$ presents a diverging peak with size increases as shown in Fig.4, which unbiasedly probes the second-order phase transition point of the model.
VI Conclusions
@@
VII Acknowledgments
–We acknowledge Xu Zhang, Zi Yang Meng, Kai Sun for useful discussions. This research was supported by the National Natural Science Foundation of China (grant nos. 12004020, 12174167, 12247101), and the Fundamental Research Funds for the Central Universities in China. The authors acknowledge Beijng PARATERA Tech CO.,Ltd. for providing HPC resources that have contributed to the research results reported within this paper. ZY thanks the start-up fund of the Westlake University.